Grade 12th passAlgebra

Simplify and express each of the following in exponential form:
(45 × a8b3 )/ (45 × a5b2 )

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

= 45 – 5 × (a8 – 5 × b3 – 2) ... [∵am ÷ an = am – n]
= 40 × (a3b)
= 1 × a3b
= a3b

​Thanks